OpenStudy (anonymous):

Solve the following quadratic equation: 12p^2 - 7p + 1 = 0

OpenStudy (anonymous):

p=1/4 or p=1/3

OpenStudy (anonymous):

Ever heard of the quadratic formula?

OpenStudy (anonymous):

\[x= \frac{ -b+-\sqrt{b ^{2}-4ac} }{ 2a }\]